# Probing Top Quark - Electron Interactions at Future Colliders

### v1

@@ II.4 Higgstrahlung and top pair production at FCC-ee
The experimental uncertainty on the top-pair production cross section measurement at the FCC-ee with $2.7~{}\textrm{ab}^{-1}$ is expected to be 0.12% [43]. The theoretical uncertainty is estimated as 1.2% by varying the top quark mass within ±$\pm 0.25$ GeV.
@@ III Numerical Fits
We are now in a position to perform a fit in which all processes are computed to NLO QCD and NLO electroweak order. We emphasize that current global fits, while accurate to NLO QCD, typically do not include NLO electroweak corrections consistently. “Partial” NLO electroweak corrections were included in the global fits of [44, 45], but the current calculation is the first to include NLO electroweak predictions for all observables, albeit for a restricted set of operators.
@@
We begin by illustrating the effects of different data sets to constrain the top quark-electron operators of eq. (2) using a $\chi^{2}$ fit. In Fig 3, we show the sensitivities of different future colliders to $C^{3311}_{qe}$ and $C^{1133}_{eu}$. All other coefficients are set to 00.
@@
Future $e^{+}e^{-}$ colliders will operate at various center-of-mass energies, ranging from the Z-pole, the $WW$ threshold ($\sqrt{s}=$ 162 GeV) and the $Zh$ threshold ($\sqrt{s}=$ 240 GeV) to the top-pair threshold ($\sqrt{s}=$ 365 GeV). The future FCC-ee and CEPC are projected to produce trillions of Z bosons at the Z-pole run, which will substantially improve the precision of Z-pole observables [42, 46]. We do a $\chi^{2}$ fit to EWPOs using the most precise theoretical results and the current experimental data, along with projections for the precision expected from FCC-ee Z pole data. (The data used in the fit is given in the supplemental material.) The current limits from EWPOs are shown in Fig. 3 as the grey band, with the improvement from FCC-ee results in the purple band.
Figure 3: The $2\sigma$ limits on $C^{3311}_{qe}$ and $C^{1133}_{eu}$ at current and future colliders. All other operators are set to 00 and $\Lambda=1$ TeV.
@@
Higgstrahlung data has a different dependence on the top-quark electron operators from the EWPOs and the limit at $\sqrt{s}=240~{}\textrm{GeV}$ yields the green band[25, 24, 45]. At $\sqrt{s}=365~{}\textrm{GeV}$, the Higgstrahlung results contribute to the fit, but the dominant contribution at this energy is from top quark pair production, which occurs at LO, shown as the dark blue band. (The results for CEPC are similar to those of FCC-ee. We do not include the ILC, as we have not optimized our study to include polarization effects relevant at the ILC.)
@@
We adopt the HL-LHC Drell-Yan projections in [47] and include the tree-level process $pp\to t\bar{t}e^{+}e^{-}$ [48] with 3 ab-1 of data. To ensure the validity of the EFT description, we restrict the Drell-Yan data to the range 500 GeV$<m_{\ell\ell}<800$ GeV. To break the degeneracies, we include both the NLO Drell-Yan and tree-level $pp\to t\bar{t}e^{-}e^{+}$ contributions, which are shown in cyan in Fig. 3. We have verified that the NLO contributions from the Drell-Yan process are numerically comparable with and complementary to the $pp\to t\bar{t}e^{-}e^{+}$ contributions, even when considering only the low-mass $m_{\ell\ell}$ bins 222 See the supplementary material for Drell-Yan results including $m_{\ell\ell}$ bins up to 4 TeV.. In the Drell-Yan contribution, the tree level process $b\bar{b}\to e^{+}e^{-}$ is also included for the operators associated with the left-handed quark doublet. The first new data after the HL-LHC will come from deep inelastic scattering at the EIC. We use 156 bins in the $x-Q^{2}$ plane with $Q^{2}>4~{}\text{GeV}^{2}$ and assume 500 fb-1 integrated luminosity for each left-handed and right-handed electron beam mode leading to the sensitivity of the orange band. The renormalization/factorization scale is $\mu=\sqrt{Q^{2}}~{}(m_{\ell\ell})$ for the DIS (Drell-Yan) process.
@@
In Fig. 3, we see that by combining different observables at a variety of energies in a global SMEFT analysis, the bounds can be significantly strengthened. However, the results from a fit to a single coefficient with the other operators marginalized over can give a quite different result from the single parameter fit, due to the presence of degeneracies. In Fig. 4, we present a marginalized fit to the operators of eq. (2) meant to demonstrate a possible sequence in time of bounds on the new physics scale $\Lambda$ (with the Wilson coefficients, $C_{i}$ set to 1). With the current EWPO measurements, there is a limit of ${\cal{O}}(\sim 2~{}{\rm{TeV}})$ for $C^{(3)}_{\ell q}$, but there are degeneracies among the remaining contributions which do not allow for a meaningful bound. Including the current bounds from the measurement of $pp\to t\bar{t}e^{+}e^{-}$ at the LHC [16] can break the degeneracies, but does not yield meaningful results as the constraints on $C_{i}$ are beyond the perturbative limit at $\Lambda=1$ TeV. Including the HL-LHC Drell–Yan and $pp\to t\bar{t}e^{+}e^{-}$ projections effectively breaks the degeneracies in the remaining four-fermion operators, as illustrated by the cyan bands in Fig. 4. The orange bands in Fig. 4 show the bounds after adding the EIC data. Finally, the results from the full FCC-ee program are shown as the red bands in Fig. 4.
@@
global-fit [-0.28,0.31] [-1.36,1.51] [-1.36,1.52] [-2.26,2.97] [-2.25,2.19]
Current Z-pole + HL-LHC+EIC single-parameter [-0.18,0.24] [-0.43,0.32] [-0.33,0.49] [-0.30,0.74] [-0.76,0.26]
global-fit [-0.27,0.25] [-1.34,1.49] [-1.33,1.38] [-2.08,2.16] [-2.13,1.62]
Current Z-pole + HL-LHC+EIC+FCC-ee single-parameter [-0.012,0.012] [-0.013,0.012] [-0.011,0.011] [-0.018,0.019] [-0.016,0.016]
global-fit [-0.059,0.061] [-0.24,0.24] [-0.21,0.21] [-0.39,0.39] [-0.34,0.33]
@@ IV Conclusions
In a global fit to the top-electron 4-fermion operators, current Z-pole observables impose a stringent constraint on the SU(2)L triplet $4$-fermion operator $\mathcal{O}^{(3),1133}_{\ell q}$, with sensitivity reaching up to $\Lambda\sim 2$ TeV, while the other four top quark-electron operators are unconstrained due to degeneracies. Including the future HL-LHC tree-level $pp\to t\bar{t}e^{+}e^{-}$ projections breaks the degeneracies, leading to weak constraints. Adding the future NLO Drell-Yan measurements will significantly improve the limits on the Wilson coefficients of the four SU(2)L singlet $4$-fermion operators by more than an order of magnitude, as shown in Table 4. Finally, the FCC-ee Z-pole run is expected to reduce the experimental uncertainties on the Z-pole observables and has the potential to push the limit on $\Lambda$ for the operator $\mathcal{O}^{(3),1133}_{\ell q}$ to 4 TeV. Nonetheless, it has very limited effects on the other four SU(2)L singlet operators due to degeneracies. The full FCC-ee program is needed to probe the electron-top $4$-fermion operators at LO and to improve the limits on $\Lambda$ for the four SU(2)L singlet operators by a factor of 2. We note that similar results can be obtained for CEPC.
@@ Acknowledgements
P.P.G. is supported by the Ramón y Cajal grant RYC2022-038517-I funded by MCIN/AEI/10.13039/501100011033 and by FSE+, and by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S. The work of LB is supported in part by the U.S. Department of Energy under Grant No. DE-SC0010102 and by the College of Arts and Sciences of Florida State University. S.D. and H.L. are supported by the U.S. Department of Energy under Grant Contract DE-SC0012704. Digital data is provided in the gitlab repositories referenced in the text.
@@
$\Gamma_{W}$(GeV) $2.085\pm 0.042$ $2.0903\pm 0.0003$[51] ±$\pm 2.7\times 10^{-4}\pm 2.0\times 10^{-4}$
$\Gamma_{Z}$(GeV) $2.4955\pm 0.0023$ $2.4943\pm{\color[rgb]{0,0,0}0.0004}~{}(8.0\times 10^{-5})$[52, 53, 54] ±$\pm 4\times 10^{-6}\pm 12\times 10^{-6}$
$R_{e}$ $20.804\pm 0.05$ $20.732\pm{\color[rgb]{0,0,0}0.006}~{}(1.2\times 10^{-3})$[52, 53, 54] ±$\pm 3.4\times 10^{-6}\pm 2.3\times 10^{-6}$
$R_{\mu}$ $20.784\pm 0.034$ $20.732\pm{\color[rgb]{0,0,0}0.006}~{}(1.2\times 10^{-3})$[52, 53, 54] ±$\pm 2.4\times 10^{-6}\pm 2.3\times 10^{-6}$
$R_{\tau}$ $20.764\pm 0.045$ $20.779\pm{\color[rgb]{0,0,0}0.006}~{}(1.2\times 10^{-3})$[52, 53, 54] ±$\pm 2.7\times 10^{-6}\pm 2.3\times 10^{-6}$
$R_{b}$ $0.21629\pm 0.00066$ $0.2159\pm 0.0001~{}(2\times 10^{-5})$[52, 53, 54] ±$\pm 1.2\times 10^{-6}\pm 1.6\times 10^{-6}$
$R_{c}$ $0.1721\pm 0.0030$ $0.1722\pm 0.00005~{}(1\times 10^{-5})$[52, 53, 54] ±$\pm 1.4\times 10^{-6}\pm 2.2\times 10^{-6}$
$\sigma_{h}$(nb) $41.481\pm 0.033$ $41.492\pm{\color[rgb]{0,0,0}0.006}~{}(5\times 10^{-4})$[52, 53, 54] ±$\pm 3\times 10^{-5}\pm 8\times 10^{-4}$
$A_{e}({\text{from}}~{}A_{LR}~{}{\text{had}})$ $0.15138\pm 0.00216$ $0.1469\pm 0.0004$[54, 55] ±$\pm 1.4\times 10^{-5}$
$A_{e}({\text{from}}~{}A_{LR}~{}{\text{lep}})$ $0.1544\pm 0.0060$ $0.1469\pm 0.0004$[54, 55]
[…]
$A_{\mu}$ $0.142\pm 0.015$ $0.1469\pm 0.0004$[54, 55] ±$\pm 3.2\times 10^{-5}$
$A_{\tau}({\text{from~{}SLD}})$ $0.136\pm 0.015$ $0.1469\pm 0.0004$[54, 55] ±$\pm 3.4\times 10^{-5}$
$A_{\tau}(\tau~{}{\text{pol}})$ $0.1439\pm 0.0043$ $0.1469\pm 0.0004$[54, 55]
$A_{c}$ $0.670\pm 0.027$ $0.66773\pm 0.0002$[54, 55] ±$\pm 6.0\times 10^{-5}$
$A_{b}$ $0.923\pm 0.020$ $0.92694\pm 0.00006$[52, 53, 54] ±$\pm 9.8\times 10^{-5}$
$A_{s}$ $0.895\pm 0.091$ $0.93563\pm 0.00004$[54, 55] ±$\pm 1.2\times 10^{-4}$
$A_{e,FB}$ $0.0145\pm 0.0025$ $0.0162\pm 0.0001$[54, 55] ±$\pm 3.3\times 10^{-6}\pm 2.4\times 10^{-6}$
$A_{\mu,FB}$ $0.0169\pm 0.0013$ $0.0162\pm 0.0001$[54, 55] ±$\pm 2.3\times 10^{-6}\pm 2.4\times 10^{-6}$
$A_{\tau,FB}$ $0.0188\pm 0.0017$ $0.0162\pm 0.0001$[54, 55] ±$\pm 2.8\times 10^{-6}\pm 2.4\times 10^{-6}$
$A_{b,FB}$ $0.0996\pm 0.0016$ $0.1021\pm 0.0003$[52, 53, 54] ±$\pm 4\times 10^{-6}\pm 4\times 10^{-6}$
$A_{c,FB}$ $0.0707\pm 0.0035$ $0.0736\pm 0.0003$[54, 55] ±$\pm 5\times 10^{-6}\pm 5\times 10^{-6}$
$A_{s,FB}$ $0.0976\pm 0.0114$ $0.10308\pm 0.0003$[54, 55] ±$\pm 7.4\times 10^{-6}\pm 7.4\times 10^{-6}$
$M_{W}$(GeV) PDG World Ave $80.377\pm 0.012$ $80.357\pm{\color[rgb]{0,0,0}0.004}~{}(0.001)$[56, 57] ±$\pm 1.8\times 10^{-4}\pm 1.6\times 10^{-4}$
@@
Table 1: The numerical values used in the Z-pole observables analysis. The experimental results are taken from the Particle Data Group[58]. The FCC-ee projections are taken from [42]. Please note that the theoretical central value of $1/\alpha_{\rm QED}$ is obtained at one-loop level, while in our analysis, we set it equal to the experimental central value.

### v2

@@ II.4 FCC-ee
The experimental uncertainty on the top-pair production cross section measurement at the FCC-ee with $2.7~\textrm{ab}^{-1}$ is expected to be 0.12% [Defranchis:2025auz]. The theoretical uncertainty is estimated as 1.2% by varying the top quark mass within $\pm 0.25$ GeV. We also include the forward-backward asymmetry of top quarks $A^{t}_{\rm FB}$, which can be calculated as
$A^{t}_{\rm FB}/10^{3}=-88.3+\biggl\{145.3(C^{(1),1133}_{\ell q}-C^{(3),1133}_{\ell q})$
$-47.4C^{1133}_{\ell u}-21.2C^{3311}_{qe}+99.9C^{1133}_{eu}\biggr\}\biggl(\frac{1~\textrm{TeV}}{\Lambda}\biggr)^{2}.$ (8)
The uncertainty is set to be 0.0016 [Lroehrig2025] including both systematic and statistical ones.
@@ III Numerical Fits
We are now in a position to perform a fit in which all processes are computed to NLO QCD and NLO electroweak order. We emphasize that current global fits, while accurate to NLO QCD, typically do not include NLO electroweak corrections consistently. “Partial” NLO electroweak corrections were included in the global fits of [Bartocci:2023nvp, Maura:2025rcv], but the current calculation is the first to include NLO electroweak predictions for all observables, albeit for a restricted set of operators.
@@
We begin by illustrating the effects of different data sets to constrain the top quark-electron operators of eq. (2) using a $\chi^{2}$ fit. In Fig 3, we show the sensitivities of different future colliders to $C^{3311}_{qe}$-$C^{1133}_{eu}$ (upper panel) and $C^{(3),1133}_{\ell q}$-$C^{(1),1133}_{\ell q}$ (lower panel). All other coefficients are set to 0.
@@
Future $e^{+}e^{-}$ colliders will operate at various center-of-mass energies, ranging from the $Z$-pole, the $WW$ threshold ($\sqrt{s}=$ 162 GeV) and the $Zh$ threshold ($\sqrt{s}=$ 240 GeV) to the top-pair threshold ($\sqrt{s}=$ 365 GeV). The future FCC-ee and CEPC are projected to produce trillions of $Z$ bosons at the $Z$-pole run, which will substantially improve the precision of $Z$-pole observables [Selvaggi:2025kmd, Maura:2024zxz]. We do a $\chi^{2}$ fit to EWPOs using the most precise theoretical results and the current experimental data, along with projections for the precision expected from FCC-ee $Z$-pole data. (The data used in the fit are given in the Supplemental Material.) The current limits from EWPOs are shown in Fig. 3 as the gray band, with the improvement from FCC-ee results in the purple band.
Figure 3: The $2\sigma$ limits on $C^{3311}_{qe}$-$C^{1133}_{eu}$ (upper panel) and $C^{(3),1133}_{\ell q}$-$C^{(1),1133}_{\ell q}$ (lower panel) at current and future colliders. In the upper panel, the FCC-ee (365 GeV) projection appears as a dot at the origin (0,0). For clarity, we mark it with a blue star. All other operators are set to 0 and $\Lambda=1$ TeV. The calculation does not include contributions of $\mathcal{O}(1/\Lambda^{4})$.
@@
Higgstrahlung data have a different dependence on the top-quark electron operators from the EWPOs and the limit at $\sqrt{s}=240~\textrm{GeV}$ yields the green band[Asteriadis:2024xts, Asteriadis:2024xuk, Maura:2025rcv]. At $\sqrt{s}=365~\textrm{GeV}$, the Higgstrahlung results contribute to the fit, but the dominant contribution at this energy is from top quark pair production, which occurs at LO, shown as the dark blue band. (The results for CEPC are similar to those of FCC-ee. We do not include the ILC, as we have not optimized our study to include polarization effects relevant at the ILC.)
@@
We adopt the HL-LHC Drell-Yan projections in [Greljo:2021kvv]. To ensure the validity of the EFT description, we restrict the Drell-Yan data to the 3 bins in the di-lepton invariant mass $m_{ee}$, which are [500,550], [550,650], and [650,800] GeV. The tree-level process $pp\to t\bar{t}e^{+}e^{-}$ is also included in our analysis. We follow [tesina_abel] using the cross section measurements with four bins in $m_{ee}$, which are [100,120], [120,140], [140,180], and >>180 GeV with 3 ab-1 of data. Including both the NLO Drell-Yan and tree-level $pp\to t\bar{t}e^{-}e^{+}$ contributions can break the degeneracies in the fit, as shown by the cyan line in Fig. 3. We have verified that the NLO contributions from the Drell-Yan process are numerically comparable with and complementary to the $pp\to t\bar{t}e^{-}e^{+}$ contributions, even when considering only the low-mass $m_{\ell\ell}$ bins  [^2] See the supplementary material for Drell-Yan results including $m_{\ell\ell}$ bins up to 4 TeV.. In the Drell-Yan contribution, the tree-level process $b\bar{b}\to e^{+}e^{-}$ is also included for the operators associated with the left-handed quark doublet. We find that the sensitivity to $C^{(3),1133}_{\ell q}/C^{(1),1133}_{\ell q}$ is dominated by the one-loop (tree-level) contribution. For $C^{(1),1133}_{\ell q}$ there is a cancellation between the up- and down-type quark contributions at the NLO, as shown in Table 2, while for $C^{3311}_{qe}$, the two contributions are comparable. The first new data after the HL-LHC will come from deep inelastic scattering at the EIC. We use 156 bins in the $x-Q^{2}$ plane with $Q^{2}>4~\text{GeV}^{2}$ and assume 500 fb-1 integrated luminosity for each left-handed and right-handed electron beam mode leading to the sensitivity of the orange band. The renormalization/factorization scale is $\mu=\sqrt{Q^{2}}~(m_{\ell\ell})$ for the DIS (Drell-Yan) process.
@@
In Fig. 3, we see that by combining different observables at a variety of energies in a global SMEFT analysis, the bounds can be significantly strengthened. However, the results from a fit to a single coefficient with the other operators marginalized over can give a quite different result from the single parameter fit, due to the presence of degeneracies. In Fig. 4, we present a marginalized fit to the operators of eq. (2) meant to demonstrate a possible sequence in time of bounds on the new physics scale $\Lambda$ (with the Wilson coefficients, $C_{i}$ set to 1). With the current EWPO measurements, there is a limit of ${\cal{O}}(\sim 2~{\rm{TeV}})$ for $C^{(3)}_{\ell q}$, but there are degeneracies among the remaining contributions which do not allow for a meaningful bound. Including the current bounds from the measurement of $pp\to t\bar{t}e^{+}e^{-}$ at the LHC [CMS:2023xyc, ATLAS:2025yww] can break the degeneracies, but does not yield meaningful results as the constraints on $C_{i}$ are beyond the perturbative limit at $\Lambda=1$ TeV. Including the HL-LHC Drell–Yan and $pp\to t\bar{t}e^{+}e^{-}$ projections effectively breaks the degeneracies in the remaining four-fermion operators, as illustrated by the cyan bands in Fig. 4. The orange bands in Fig. 4 show the bounds after adding the EIC data. Finally, the results from the full FCC-ee program are shown as the red bands in Fig. 4.
@@
global-fit [-0.28,0.31] [-1.37,1.53] [-1.37,1.39] [-2.29,3.00] [-2.27,2.21]
Current $Z$-pole + HL-LHC+EIC single-parameter [-0.17,0.23] [-0.36,0.28] [-0.30,0.43] [-0.28,0.57] [-0.66,0.26]
global-fit [-0.27,0.24] [-1.34,1.46] [-1.33,1.38] [-2.02,1.88] [-2.10,1.46]
Current $Z$-pole + HL-LHC+EIC
+FCC-$e$
single-parameter [-0.0012,0.0012] [-0.0012,0.0012] [-0.0078,0.0078] [-0.0033,0.0034] [-0.016,0.015]
global-fit [-0.058,0.059] [-0.029,0.030] [-0.070,0.065] [-0.12,0.13] [-0.11,0.11]
@@ IV Conclusions
In a global fit to the top-electron 4-fermion operators, current $Z$-pole observables impose a stringent constraint on the SU(2)L triplet $4$-fermion operator $\mathcal{O}^{(3),1133}_{\ell q}$, with sensitivity reaching up to $\Lambda\sim 2$ TeV, while the other four top quark-electron operators are unconstrained due to degeneracies. Including the future HL-LHC tree-level $pp\to t\bar{t}e^{+}e^{-}$ projections breaks the degeneracies, leading to weak constraints. Adding the future NLO Drell-Yan measurements will significantly improve the limits on the Wilson coefficients of the four SU(2)L singlet $4$-fermion operators by more than an order of magnitude, as shown in Table 3. Finally, the FCC-ee $Z$-pole run is expected to reduce the experimental uncertainties on the $Z$-pole observables and has the potential to push the limit on $\Lambda$ for the operator $\mathcal{O}^{(3),1133}_{\ell q}$ to 4 TeV. Nonetheless, it has very limited effects on the other four SU(2)L singlet operators due to degeneracies. The full FCC-ee program is needed to probe the electron-top $4$-fermion operators at LO and to improve the limits on $\Lambda$ for the four SU(2)L singlet operators by a factor of $\sim 3-6$. We note that similar results can be obtained for CEPC.
@@ Acknowledgements
We thank Admir Greljo for pointing out the importance of the $R_{b}$ measurement above the $Z$-pole at the FCC-ee. P.P.G. is supported by the Ramón y Cajal grant RYC2022-038517-I funded by MCIN/AEI/10.13039/501100011033 and by FSE+, and by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S. The work of LB is supported in part by the U.S. Department of Energy under Grant No. DE-SC0010102 and by the College of Arts and Sciences of Florida State University. S.D. and H.L. are supported by the U.S. Department of Energy under Grant Contract DE-SC0012704. Digital data is provided in the supplemental files.
Probing Top Quark Interactions at Future Colliders
@@
$\Gamma_{W}$(GeV) $2.085\pm 0.042$ $2.0903\pm 0.0003$[Cho:2011rk] $\pm 2.7\times 10^{-4}\pm 2.0\times 10^{-4}$
$\Gamma_{Z}$(GeV) $2.4955\pm 0.0023$ $2.4943\pm{\color[rgb]{0,0,0}0.0004}~(8.0\times 10^{-5})$[Freitas:2014hra, Dubovyk:2018rlg, Dubovyk:2019szj] $\pm 4\times 10^{-6}\pm 12\times 10^{-6}$
$R_{e}$ $20.804\pm 0.05$ $20.732\pm{\color[rgb]{0,0,0}0.006}~(1.2\times 10^{-3})$[Freitas:2014hra, Dubovyk:2018rlg, Dubovyk:2019szj] $\pm 3.4\times 10^{-6}\pm 2.3\times 10^{-6}$
$R_{\mu}$ $20.784\pm 0.034$ $20.732\pm{\color[rgb]{0,0,0}0.006}~(1.2\times 10^{-3})$[Freitas:2014hra, Dubovyk:2018rlg, Dubovyk:2019szj] $\pm 2.4\times 10^{-6}\pm 2.3\times 10^{-6}$
$R_{\tau}$ $20.764\pm 0.045$ $20.779\pm{\color[rgb]{0,0,0}0.006}~(1.2\times 10^{-3})$[Freitas:2014hra, Dubovyk:2018rlg, Dubovyk:2019szj] $\pm 2.7\times 10^{-6}\pm 2.3\times 10^{-6}$
$R_{b}$ $0.21629\pm 0.00066$ $0.2159\pm 0.0001~(2\times 10^{-5})$[Freitas:2014hra, Dubovyk:2018rlg, Dubovyk:2019szj] $\pm 1.2\times 10^{-6}\pm 1.6\times 10^{-6}$
$R_{c}$ $0.1721\pm 0.0030$ $0.1722\pm 0.00005~(1\times 10^{-5})$[Freitas:2014hra, Dubovyk:2018rlg, Dubovyk:2019szj] $\pm 1.4\times 10^{-6}\pm 2.2\times 10^{-6}$
$\sigma_{h}$(nb) $41.481\pm 0.033$ $41.492\pm{\color[rgb]{0,0,0}0.006}~(5\times 10^{-4})$[Freitas:2014hra, Dubovyk:2018rlg, Dubovyk:2019szj] $\pm 3\times 10^{-5}\pm 8\times 10^{-4}$
$A_{e}({\text{from}}~A_{LR}~{\text{had}})$ $0.15138\pm 0.00216$ $0.1469\pm 0.0004$[Dubovyk:2019szj, Awramik:2006uz] $\pm 1.4\times 10^{-5}$
$A_{e}({\text{from}}~A_{LR}~{\text{lep}})$ $0.1544\pm 0.0060$ $0.1469\pm 0.0004$[Dubovyk:2019szj, Awramik:2006uz]
$A_{e}({\text{from~Bhabba~pol}})$ $0.1498\pm 0.0049$ $0.1469\pm 0.0004$[Dubovyk:2019szj, Awramik:2006uz]
$A_{\mu}$ $0.142\pm 0.015$ $0.1469\pm 0.0004$[Dubovyk:2019szj, Awramik:2006uz] $\pm 3.2\times 10^{-5}$
$A_{\tau}({\text{from~SLD}})$ $0.136\pm 0.015$ $0.1469\pm 0.0004$[Dubovyk:2019szj, Awramik:2006uz] $\pm 3.4\times 10^{-5}$
$A_{\tau}(\tau~{\text{pol}})$ $0.1439\pm 0.0043$ $0.1469\pm 0.0004$[Dubovyk:2019szj, Awramik:2006uz]
$A_{c}$ $0.670\pm 0.027$ $0.66773\pm 0.0002$[Dubovyk:2019szj, Awramik:2006uz] $\pm 6.0\times 10^{-5}$
$A_{b}$ $0.923\pm 0.020$ $0.92694\pm 0.00006$[Freitas:2014hra, Dubovyk:2018rlg, Dubovyk:2019szj] $\pm 9.8\times 10^{-5}$
$A_{s}$ $0.895\pm 0.091$ $0.93563\pm 0.00004$[Dubovyk:2019szj, Awramik:2006uz] $\pm 1.2\times 10^{-4}$
$A_{e,FB}$ $0145\pm 0.0025 0.0$$0162\pm 0.0001$[Dubovyk:2019szj, Awramik:2006uz] $\pm 3.3\times 10^{-6}\pm 2.4\times 10^{-6}$
$A_{\mu,FB}$ $0169\pm 0.0013 0.0$$0162\pm 0.0001$[Dubovyk:2019szj, Awramik:2006uz] $\pm 2.3\times 10^{-6}\pm 2.4\times 10^{-6}$
$A_{\tau,FB}$ $0188\pm 0.0017 0.0$$0162\pm 0.0001$[Dubovyk:2019szj, Awramik:2006uz] $\pm 2.8\times 10^{-6}\pm 2.4\times 10^{-6}$
$A_{b,FB}$ $0.0996\pm 0.0016$ $0.1021\pm 0.0003$[Freitas:2014hra, Dubovyk:2018rlg, Dubovyk:2019szj] $\pm 4\times 10^{-6}\pm 4\times 10^{-6}$
$A_{c,FB}$ $0707\pm 0.0035 0.0$$0736\pm 0.0003$[Dubovyk:2019szj, Awramik:2006uz] $\pm 5\times 10^{-6}\pm 5\times 10^{-6}$
$A_{s,FB}$ $0.0976\pm 0.0114$ $0.10308\pm 0.0003$[Dubovyk:2019szj, Awramik:2006uz] $\pm 7.4\times 10^{-6}\pm 7.4\times 10^{-6}$
$M_{W}$(GeV) PDG World Ave $80.377\pm 0.012$ $80.357\pm{\color[rgb]{0,0,0}0.004}~(0.001)$[Awramik:2003rn, Erler:2019hds] $\pm 1.8\times 10^{-4}\pm 1.6\times 10^{-4}$
@@
Table 1: The numerical values used in the $Z$-pole observables analysis. The experimental results are taken from the Particle Data Group[ParticleDataGroup:2018ovx]. The FCC-ee projections are taken from [Selvaggi:2025kmd]. Please note that the theoretical central value of $1/\alpha_{\rm QED}$ is obtained at one-loop level, while in our analysis, we set it equal to the experimental central value.